Let X and Y be independent random variables with pdf 2-y , 0sys2 2 f(x) 0, otherwise 0, otherwise ) Find E(XY) b) Find Var (2X+3Y)
Answer #1

f(x)=rac{x}{2} , 0leq x leq 2

E(x)=int_{0}^{2} xrac{x}{2} dx

=4/3

2 0

=2

V (x)=E(x^2)-[E(x)]^2

=2-[rac{4}{3}]^2

=2/9

2-y

E(y)=int_{0}^{2}yrac{2-y}{2}dy

=2/3

2-y

=2/3

V(y)=2/3-(2/3)^2

=2/9

Here x and y are independent

So,

E(xy)=E(x)E(y)

=(4/3)*(2/3)

=8/9

V(2x+3y)= 2^2V(x)+3^2V(y)

=4V(x)+9V(y)

=4*2/9 +9*2/9

=26/9